In two H-atoms their electrons are moving in 1st state and in 1st excited state respectively, ratio of time periods will be
1 : 8
1 : 4
1 : 2
2 : 1








Grade:11

1 Answers

vikas askiitian expert
13 years ago

time period =T=2pir/v                       ( r is radius of orbit,v is velocity of elctron)

 

v is inversly proportional to n

r is directly proportional to n2

 so time period (T) is directly proportionla to n3 ....

  for  1 state n =1  , for first exited state n=2

    T1/T2= (n1/n2)3

      T1/T2= 1/8     

first option is correct